THE ROTATIONAL SPECTRUM AND MOLECULAR PROPERTIES OF CHLORYL CHLORIDE, ClClO_{2}

\underline{HOLGER S. P. MÜLLER}, AND EDWARD A. COHEN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109.

ClClO_{2} is slightly higher in energy than its structural isomer ClOOCl,^{a} which has been implicated in polar ozone depletion processes. ClClO_{2} may be formed from other Cl_{x}O_{y} compounds on or in stratospheric ices.^{b} ^{c} Therefore, its millimeter and submillimeter spectrum has been studied in a flowing system as a product of the FClO_{2} + HCl reaction. The pyramidal ClClO_{2} is an asymmetric prolate top, κ = -0.7598 for ^{35}Cl^{35}ClO_{2}. It has C_{s} symmetry with a strong dipole component along the c-axis and a smaller one along the a-axis. The highest quantum numbers accessed are larger than 50 and 30 for J and K_{a}, respectively, permitting rotational and centrifugal distortion constants to be determined precisely. Splittings due to both Cl nuclei have been resolved, and a quadrupole analysis will be presented. The molecular structure has been derived from isotopomers involving ^{35}Cl and ^{37}Cl. The results will be compared with those from an earlier matrix-isolation study,^{d} from ab initio calculations,^{a} and from data of related molecules.

aT. J. Lee, C. McMichael Rohlfing, and J. E. Rice, J. Chem. Phys., 97, 6593, (1992)

bC. J. Pursell, J. Conyers, and C. Denison, J. Phys. Chem., 100, 15450, (1996)

cJ. D. Graham, J. T. Roberts, L. D. Anderson, and V. H. Grassian, J. Phys. Chem. 100, 19551, (1996)

dH. S. P. Müller and H. Willner, Inorg. Chem. 31, 2527, (1992)